If (3x-4y)+i(x+y)=7 {find x,y}
Answer by ikleyn(52879)

.
If (3x-4y)+i(x+y)=7 {find x,y}
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

You forgot to point in the condition that x and y are real numbers.
So, I will assume it is so.

If (3x-4y) + i(x+y) = 7 then

3x - 4y = 7,   (1)     (real part is 7)
 x +  y = 0.   (2)     (imaginary part is 0).

From (2), x = -y. Substitute it into (1). You will get

3*(-y) - 4y = 7,  or

-7y = 7  --->  y = -1.

So, you just found y = -1. 

Then x = -y = -(-1) = 1.

Answer. x = 1, y = -1.

There is a bunch of my lessons on complex numbers
    - Complex numbers and arithmetical operations on them
    - Complex plane
    - Addition and subtraction of complex numbers in complex plane
    - Multiplication and division of complex numbers in complex plane

    - Solved problems on arithmetic operations on complex numbers
    - Miscellaneous problems on complex numbers
in this site.

Also, you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook.